OpenStudy (muddahamburglar):

whats the period of the function f(x)= 4cot(( pi x))/6

OpenStudy (irishboy123):

is that \(f(x)= 4\dfrac {\cot \pi x}{6}\) or \(f(x)= 4 \cot\dfrac { \pi x}{6}\) use the drawing tool if you like, but suspect the latter generally, trig functions have period *T* when you have, as say in this case, \(y = 4 \cot \dfrac{2 \pi }{T}x\) so that as as x goes from 0 to \(T\), \( \dfrac{2 \pi x}{T}\) goes from 0 to \(2 \pi\), if that makes sense? if you get that, you can pattern-match out the answer already.